Circular No. 7032

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998ef IN UGC 646 W. D. Li, M. Modjaz, E. Halderson, T. Shefler, J. Y. King, M. Pa-

penkova, R. R. Treffers, and A. V. Filippenko, University of California at Berkeley, report their discovery of an apparent supernova during the course of the Lick Observatory Supernova Search with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 1998ef was discovered on an unfiltered image taken on Oct. 18.3 UT (mag ∼ 15.2), and confirmed on an earlier image taken on Oct. 14.3 (mag ∼ 16.7). SN 1998ef is located at α = 1^{h}03^{m}26^{s}.87, δ = +32^{o}14′12′′.4 (equinox 2000.0), which is ∼ 6′′.1 east and 2′′.1 south of the nucleus of UGC 646. A KAIT image of the same field on Oct. 10.3 showed nothing at the position of the new star (limiting mag ∼ 18.0). Filippenko further reports that inspection of a CCD spectrum of SN 1998ef obtained on Oct. 19 by C. De Breuck (Lawrence Livermore National Laboratory) with the 3-m Shane reflector at Lick reveals that the object is a type-Ia supernova, with well-developed Si II lines.

XTE J1906+09 T. Takeshima, Goddard Space Flight Center and Universities Space

Research Association; and T. Murakami, Institute of Space and Astronautical Science, on behalf of the RXTE Science Operations Center and ASCA team, report: “Simultaneous observations of SGR 1900+14 with ASCA (Sept. 16.794–17.831 UT) and RXTE/PCA (Sept. 16.826–16.988 and 17.281–17.388) confirmed that the 89-s pulsar XTE J1906+09 (Marsden et al. 1998, Ap.J. 502, L129) is a separate source from SGR 1900+14 (IAUC 7008). While RXTE/PCA detected significant 89-s pulsations, ASCA saw only 5.17-s pulsations from SGR 1900+14. Scans with the RXTE/PCA on Sept. 21.45 localized the 89-s pulsar to α = 19^{h}05^{m}20^{s}, δ = +9^{o}02′.5 (equinox 2000.0), which was ∼ 2′ outside the ASCA/GIS field-of-view. The error radius of 2′.0 (90-percent confidence level) is dominated by systematic uncertainties due to intensity variations of the 89-s pulsar and of nearby sources during the scans. The average flux of XTE J1906+09 and its pulse fraction during the scan observation were 4.1 mCrab (2–10 keV) and 66 percent, respectively. The pulsating x-ray flux declined below 0.3 mCrab (2–10 keV) by Oct. 2, but observations at other wavelengths to identify the counterpart are encouraged.”

1998 October 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green